# Collision-Free Robot Navigation in Crowded Environments using Learning based Convex Model Predictive Control

### 03 Mar 2024 v1

@@ Abstract
The advent of deep reinforcement learning (DRL) has significantly expanded the application range for autonomous robots. However, safe navigation in crowded and complex environments remains a persistent challenge. This study proposes a robot navigation strategy that utilizes DRL, conceptualizing the observation as the convex static obstacle-free region, a departure from traditional reliance on raw sensor inputs. The novelty of this work is threefold: (1) Formulating an action space that includes both short-term and long-term reference points, based on the robot’s kinematic limits and the convex region computed from 2D LiDAR sensor data. (2) Exploring a hybrid solution that combines DRL with Model Predictive Control (MPC). (3) Designing a customized state space and reward function based on the static obstacle-free region, reference points, and the trajectory optimized by MPC. The effectiveness of these improvements has been confirmed through experimental results, demonstrating improved navigation performance in crowded and complex environments.
@@ I INTRODUCTION
Robot navigation in densely populated environments is still a challenge that has drawn significant attention from the global research community. This interest is heightened by the integration of advanced learning methodologies. The unpredictable dynamics of these environments necessitate a robust and efficient autonomous navigation strategy. Recent advances in machine learning and deep reinforcement learning (DRL) have propelled the exploration of neural networks for navigation in these challenging settings [1].
Traditional DRL-based navigation methods, which rely on training robots in simulated environment using simulated sensory inputs, often face real-world limitations due to discrepancies between simulation and reality, particularly with regard to noise and friction [2]. A significant challenge is the dependency on neural networks to issue action commands, a factor that can diminish performance when shifting from simulated environments to actual ones [3]. To address this challenge, researchers, including Gao et al. [4], have introduced innovative strategies. One effective approach involves separating the control mechanism from the neural network and employing control methodologies that are viable in real-world settings.
Inspired by these advancements, our study employs Model Predictive Control (MPC), valued for its ability to enforce hard constraints, ensuring control commands stay within predefined limits [5]. Consequently, we define the network output as a series of reference points for MPC’s reference trajectory. Furthermore, we utilize the convex static obstacle-free region derived from LiDAR point clouds as the constraint space. This strategy confines the MPC-optimized trajectory strictly within the convex region, thereby ensuring navigation that is both kinematically feasible and secure.
@@
  • Designing the action and action space based on the convex static obstacle-free region derived from LiDAR data. To facilitate both short-term dynamic obstacle avoidance and long-term navigation, we characterize the action as the reference trajectory composed of reference points at specific time intervals. The action space is defined as the intersection of the convex region with the robot’s motion limits, from which we sample static safe reference trajectory. Subsequently, MPC is employed to track the reference trajectory, incorporating the convex region constraint within the MPC formulation to ensure the generation of safe and reliable control commands. Furthermore, we develop customized state space and reward function based on the convex region, reference points, and MPC-optimized trajectory.
  • Implementing a seven stage curriculum training strategy, tested through a series of detailed experiments. These included ablation studies that analyze effects of various action space designs and reward function configurations on navigation performance. Furthermore, a comparative evaluation against the Timed Elastic Band (TEB) method highlighted our training approach’s superior performance.
  • We have made our work open source by sharing both the implementation and benchmark data in a code repository accessible at https://github.com/sunnyhello369/flat_ped_sim.git.
@@ II RELATED WORK
Traditional navigation approaches for robots in crowded spaces, such as the velocity-based, force-based, and potential field methods [6], face challenges due to their reliance on simplistic assumptions about pedestrian dynamics. The Social Force Model (SFM) [7], along with techniques like Reciprocal Velocity Obstacles (RVO) [8] and ORCA [9], provide foundational frameworks for predicting pedestrian movements, yet they may not fully capture the unpredictability of real-world environments [10].
Chen et al. [11] have advanced the field by introducing an interactive Model Predictive Control (iMPC) framework that utilizes the iORCA model for enhanced prediction of pedestrian movements, thereby improving robot navigation in crowded environments.
DRL methods are also applied for crowd navigation. Yao et al. [12] developed an end-to-end DRL network that, by distinguishing between obstacles and pedestrians and utilizing sensor data along with pedestrian maps, significantly enhances robotic navigation capability. On a related front, Wang et al. [13] crafted a hierarchical network of gated-recurrent-unit (GRU) for environment modeling to predict dynamic changes, which focuses on dynamic environment representation, adaptive perception mechanism, and adaptive reward function for collision-free motion planning policy.
@@
In parallel, hybrid methods have also seen novel applications. Gao et al. [4] developed a method, encoding vehicle and predicted trajectories into binary image as the state of DRL framework. Action is defined as the vehicle’s future positions in polar coordinates. They applied convex optimization to these discrete reference points, transforming them into a reachable trajectory for the vehicle to follow. This method decouples control from other processes, with DRL concentrating on suggesting reference positions and convex optimization ensuring the trajectory’s practicality for vehicle implementation.
@@
Linh et al. [14] have explored various planning methods, including traditional MPC, TEB, and DRL-based end-to-end navigation strategies. They designed a policy network as a navigation method selector, choosing the appropriate navigation strategy based on static and dynamic environmental information. Brito et al. [3] have proposed a DRL-based decision-planning method that recommends target positions for the MPC planner, facilitating navigation that considers interactions with other agents.
@@
Our navigation strategy is similar to that of Gao et al. [4], where the action is defined as a sequence of future positions for the robot. However, to achieve better navigation performance and smoother motion trajectories in dynamic, crowded settings, our method employs a state observation based on the convex static obstacle-free region calculated from 2D LiDAR data. Furthermore, we incorporate the convex region into the design of the action-state space, the reward function, and our MPC framework used for tracking the reference trajectory.
@@ III-A Problem Definition
We model our problem using a Partially Observable Markov Decision Process (POMDP), which is well-suited for environments with inherent uncertainties and dynamic conditions. Our POMDP model, $M=(S,A,T,R,\gamma)$ , includes: state space S, action space A, state transition function T, reward function R, and discount factor $\gamma\in[0,1)$. At time t, the action of the agent, $a_{t}$, consists of the short-term and long-term reference points, $Q_{t}^{s}$ and $Q_{t}^{l}$, is formulated in Section III-C and the reward $r_{t}$ detailed in Section III-F. The state of the agent, $s_{t}$, is defined in Section III-E.
@@ III-B Agent Dynamics
To facilitate computation in planning and control, the omni-directional mobile robot is simplified to a point mass. This simplification turns the robot’s kinematics into a 2D third-order integral model. The model’s state, $x=\left[p_{x},p_{y},v_{x},v_{y},a_{x},a_{y}\right]^{T}$, captures the position, velocity, and acceleration, while the control input, $u=\left[j_{x},j_{y}\right]^{T}$, denotes the jerk.
@@ III-C Action Space Design Based on Convex Region
The convex static obstacle-free region is generated by the algorithm proposed by Zhong et al. [15], which can efficiently create reliable convex spaces among obstacles of any shape from LiDAR point cloud data.
To navigate around dynamic obstacles effectively, our action is designed as a sequence of reference points at specific time intervals. To reduce training complexity, this sequence is simplified to two points: one representing the robot’s reference position after a control cycle ($t_{c}$) and the other after an MPC prediction horizon ($T$), corresponding to short-term and long-term reference points, respectively. The path from the robot’s current position to these points forms the reference trajectory for MPC tracking.
In the design of the action space, we firstly ensure that the sampling of reference points is confined within the convex region. Then, centering on the robot’s current position $O$ and considering its kinematic limits, we define two circular areas. These circles, intersected with the convex region, constitute our short-term and long-term action spaces, as illustrated by the red areas in Fig. 1 and 2. The policy network is designed to output two pairs of two-dimensional data, which, after sigmoid activation, yield $\hat{a}_{t}=\left\{\left(\alpha_{t}^{s},\beta_{t}^{s}\right),\left(\alpha_{t}^{l},\beta_{t}^{l}\right)|\alpha_{t},\beta_{t}\in\left(0,1\right)\right\}$.
Figure 1: Schematic of Short-Term Reference Point Formulation
The short-term reference point’s coordinate, $Q_{t}^{s}$, is calculated as Equation 2, with Fig. 1 illustrating this process. A ray at angle $\theta_{t}^{s}$ intersects with the short-term action space at $P_{t}^{s}$. We determine $P_{t}^{s}$ by calculating the polar angles of the convex’s vertices and finding which edge the angle $\theta_{t}^{s}$ falls within. This intersection, is solved by combining the ray and the edge equations, yields $P_{t}^{s}$ in the LiDAR Cartesian system. Multiplying the distance from $O$ to $P_{t}^{s}$, $d_{\theta,t}^{s}$, by the distance scaling factor $\beta_{t}^{s}$, we get the distance $d_{t}^{s}$, as formulated in Equation 2.
@@
Figure 2: Schematic of Long-Term Reference Point Formulation
The coordinate of the long-term reference point, $Q_{t}^{l}$ , relative to $O$, is calculated as Equation 3, with Fig. 2 illustrating the process. We establish a new polar coordinate system centered at the short-term reference point $Q_{t}^{s}$. The long-term reference angle, $\theta_{t}^{l}$, is calculated by adding a increment angle, $\alpha_{t}^{l}\cdot 2\pi$ , to $\theta_{t}^{s}$. A ray at angle $\theta_{t}^{l}$ intersects with the long-term action space at $P_{t}^{l}$. Following the procedure used for the short-term position, we can determine the coordinates of $Q_{t}^{l}$.
$\theta_{t}^{l}$ $=\theta_{t}^{s}+\alpha_{t}^{l}\cdot 2\pi$ (3)
@@
To navigate safely in unknown environment, our method leverages real-time LiDAR data to generate the convex static obstacle-free region for local navigation. By ensuring the reference trajectory remains within the lasted convex region, collision with static obstacles can be avoided. The reference trajectory, connecting both long-term and short-term reference points, illustrated by the green solid line in Fig. 3. Subsequently, the MPC framework solves the local navigation problem by computing the optimal control inputs, thereby enabling the robot to adhere to the reference trajectory. This process is iteratively repeated, ensuring the robot can follow updated trajectory and reach its target without colliding with obstacles.
@@ III-D Model Predictive Control Formulation
Upon the discretization of the two-dimensional, third-order integral model governing robotic dynamics (referenced in Section III-B, Equation 1), the sequential relationship is obtained as $x_{i}=Fx_{i-1}+Gu_{i-1}$ for $i=\{1,\ldots,N\}$. In this formulation, $F$ signifies the matrix defining the discrete-time state transitions, while $G$ delineates the matrix characterizing the discrete-time control inputs.
Using the method mentioned in Section III-C, we get a convex region from point cloud data, with its vertices $P_{j}^{c},j=\left\{1,...,rnum_{v}\right\}$ arranged in clockwise direction. This region limits the MPC-optimized points to ensure safety by preventing collisions with static obstacles. This constraint is efficiently met by determining whether the trajectory point $Q_{i}$ is inside the convex region through a streamlined vector method $\overrightarrow{Q_{i}P_{j}^{c}}\times\overrightarrow{Q_{i}P_{j+1}^{c}}\leq 0$.
During navigation, this vector method also is used to verify if the goal endpoint lies within the convex region. If so, this endpoint is directly used as the long-term reference point $Q_{t}^{l}$ , which promotes early training by facilitating the acquisition of high-reward trajectories. And we also introduce a final state stop cost, $\cos t_{stop}$, into the MPC’s cost function when the goal is within the convex region. This cost aims to minimize both the velocity $V_{N}=\left(v_{x}^{N},v_{y}^{N}\right)$ and acceleration $A_{N}=\left(a_{x}^{N},a_{y}^{N}\right)$ at the destination, ensuring a controlled and safe termination of movement. The final state stop cost $\cos t_{stop}$ is defined in Equation 4.
The short-term and long-term reference points are used to guide the positions of the first and last control period states predicted by the MPC. The cost function incorporates discrepancies between MPC-predicted points and reference points to ensure fidelity to the reference trajectory. Ultimately, the MPC formulation is approached as a quadratic programming problem, formulated in Equation 5, optimizing the control sequence $u_{0:N-1}^{*}$ The optimization produces a series of MPC-optimized points $Q_{i}^{*},i=\left\{1,...,N\right\}$, forming the MPC-optimized trajectory, as depicted in Fig. 4.
Figure 4: Schematic of MPC Optimization Process
$\cos t_{stop}=\begin{cases}\begin{aligned} &w_{vend}\|V_{N}\|^{2}\\ &+w_{aend}\|A_{N}\|^{2}\end{aligned}&\text{if goal in convex}\\ 0&\text{otherwise}\end{cases}$ (4)
@@
$\quad+w_{smooth}\sum_{k=0}^{N-1}{\left\|u_{k}\right\|^{2}}+\cos t_{stop}$

### 14 Mar 2024 v2

@@ Abstract
Navigating robots safely and efficiently in crowded and complex environments remains a significant challenge. However, due to the dynamic and intricate nature of these settings, planning efficient and collision-free paths for robots to track is particularly difficult. In this paper, we uniquely bridge the robot’s perception, decision-making and control processes by utilizing the convex obstacle-free region computed from 2D LiDAR data. The overall pipeline is threefold: (1) We proposes a robot navigation framework that utilizes deep reinforcement learning (DRL), conceptualizing the observation as the convex obstacle-free region, a departure from general reliance on raw sensor inputs. (2) We design the action space, derived from the intersection of the robot’s kinematic limits and the convex region, to enable efficient sampling of inherently collision-free reference points. These actions assists in guiding the robot to move towards the goal and interact with other obstacles during navigation. (3) We employ model predictive control (MPC) to track the trajectory formed by the reference points while satisfying constraints imposed by the convex obstacle-free region and the robot’s kinodynamic limits. The effectiveness of proposed improvements has been validated through two sets of ablation studies and a comparative experiment against the Timed Elastic Band (TEB), demonstrating improved navigation performance in crowded and complex environments.
@@ I INTRODUCTION
Robot navigation in crowded environments is still a challenge that has drawn significant attention from the global research community [1]. This interest is heightened by the integration of advanced learning methodologies [2], [3]. A reliable and effective autonomous navigation strategy is required due to the unpredictable dynamics. Recent advances in machine learning and DRL have facilitated the investigation of neural networks for navigation in these challenging settings [4], [5], [6].
Current DRL-based end to end navigation strategies typically define the expected speed, acceleration, and other control variables in different dimensions as actions, within continuous action spaces [7], [8]. Alternatively, these control variables are integrated into predefined actions (e.g., moving forward, braking) to form discrete action spaces [9], [10]. Nevertheless, neither continuous nor discrete actions directly depict the robot’s expected trajectory. Akmandor et al. [11] proposed mapping control variables to a set of motion primitives, enabling the agent to select the optimal trajectory to track. Nevertheless, these methods still cannot guarantee the interpretability and safety of the robot’s motion, losing explicit mechanism for the integration of optimization/decision and control, thereby leading to uninterpretable autonomy and impractical implementation. To address this challenge, researchers, including Gao et al. [12], have introduced innovative approaches. One effective approach is to separate the control mechanism from the neural network and employ control methods that are feasible in real-world settings.
Figure 1: Proposed Navigation Architecture: The convex obstacle-free region is obtained from 2D LiDAR point cloud data. The policy network selects reference points ($Q_{t}^{s}$ and $Q_{t}^{l}$) from this convex region based on consecutive frames of observations. The MPC is then employed with online optimization to compute optimal local trajectory. The trajectory follows the reference points closely while satisfying both kinodynamic constraints and convex obstacle-free region requirements. This iterative process continues until the robot reaches its goal.
Inspired by these advancements, our study employs MPC, valued for its ability to enforce hard constraints, ensuring control commands stay within predefined limits [13]. Consequently, we define the action as a sequence of reference points for MPC’s reference trajectory. Furthermore, we utilize the convex obstacle-free region derived from LiDAR point clouds as the constraint space, as depicted in Fig. 1. This strategy confines the MPC-optimized trajectory strictly within the convex region, thereby ensuring navigation that is both kinematically feasible and safe.
@@
  • We design the action and action space based on the convex obstacle-free region derived from LiDAR data. To facilitate both short-term dynamic obstacle avoidance and long-term navigation, we characterize the action as a reference trajectory composed of reference points at specific time intervals. The action space is defined as the intersection of the convex region with the robot’s motion limits, from which we sample collision-free reference points efficiently. Subsequently, MPC is employed to track the trajectory, incorporating the convex region constraint within the MPC formulation to ensure the generation of safe and reliable control commands. Furthermore, we develop customized state space and reward function based on the convex region, reference points, and MPC-optimized trajectory.
  • We implement a seven-stage curriculum training strategy and evaluate it through detailed experiments. The experiments include ablation studies to analyze the impact of varying action spaces and reward functions on navigation performance. Furthermore, a comparative evaluation with TEB method highlighted the superior performance of our method.
  • We have made our work open source by publishing both the implementation and benchmark data in a code repository accessible at https://github.com/sunnyhello369/flat_ped_sim.git.
@@ II RELATED WORK
Previous robot navigation approaches often consider velocity, force, potential, and other factors as key concepts [14]. However, they face challenges due to their reliance on simplistic assumptions about pedestrian dynamics. The Social Force Model (SFM) [15], along with methods like Reciprocal Velocity Obstacles (RVO) [16] and Optimal Reciprocal Collision Avoidance (ORCA) [17], provide foundational frameworks for predicting pedestrian movements, yet they may not fully capture the unpredictability of real-world environments [18]. Chen et al. [19] proposed an interactive Model Predictive Control (iMPC) framework that utilizes the iORCA model for enhanced prediction of pedestrian movements, thereby improving robot navigation in crowded environments.
Learning based methods are also applied to navigation in crowded environments. Yao et al. [20] developed an end-to-end DRL framework that uses sensor data and pedestrian maps to distinguish between obstacles and pedestrians to enhance dynamic obstacle avoidance capabilities. Hu et al. [21] presented a planning-oriented framework that utilizes Transformer-based models [22] for perception, prediction, and planning tasks. This framework overcomes the limitations of modular designs and multi-task learning, achieving end-to-end navigation.
@@
In parallel, hybrid methods have also seen novel applications. Gao et al. [12] developed a method, encoding vehicle and predicted trajectories into binary images as the state. Action is defined as vehicle’s future positions in polar coordinate. They applied convex optimization to these discrete reference points, transforming them into reachable trajectory to track. This method decouples control from other processes, with DRL focuses on suggesting reference positions and convex optimization ensuring the trajectory’s feasibility for vehicle execution.
@@
Nikdel et al. [23] introduced a hybrid method that combines DRL with classical trajectory planning, where DRL estimates human trajectories and suggests short-term goals. These goals are then used by TEB to navigate the robot in front of humans. Linh et al. [24] had explored various planning methods, including MPC, TEB, and DRL-based end-to-end navigation strategies. They designed policy network as a strategy selector, choosing the appropriate strategy based on static and dynamic environmental information. Brito et al. [25] proposed a DRL-based decision-planning method that recommends target positions for the MPC planner. Facilitate navigation by considering interactions with other agents. Compared to [25], our method does not require states of obstacles to be observed, only requires raw 2D LiDAR point cloud data.
@@
Our navigation strategy is similar to that of Gao et al. [12], where the action is defined as a sequence of future positions for the robot. However, to achieve better navigation performance and smoother motion trajectories in dynamic crowded environments, our method utilizes the state observation based on the convex obstacle-free region. Furthermore, we integrate the convex region into the architecture of the action-state space, the reward function, and our MPC framework.
@@ III-A Problem Definition
We model our problem using a Partially Observable Markov Decision Process (POMDP), which is well-suited for environments with inherent uncertainties and dynamic conditions. The POMDP model $M=(S,A,T,R,\Omega,O,\gamma)$ consists of state space $S$, action space $A$, state transition function $T$, reward function $R$, finite set of observations $\Omega$, observation function $O$, and discount factor $\gamma\in[0,1)$. At time $t$, the action $a_{t}$, consists of the short-term and long-term reference points, $Q_{t}^{s}$ and $Q_{t}^{l}$, is formulated in Section III-C and the reward $r_{t}$ detailed in Section III-F. The state $s_{t}$ is defined in Section III-E.
@@ III-B Agent Dynamics
To facilitate computation in planning and control, the omni-directional mobile robot is simplified to a point mass. This simplification converts the robot’s kinematics into a 2D third-order integral model. The state of the model denoted by $x=\left[p_{x},p_{y},v_{x},v_{y},a_{x},a_{y}\right]^{T}$, represents the position, velocity, and acceleration. On the other hand, the control input, $u=\left[j_{x},j_{y}\right]^{T}$, denotes the jerk.
@@ III-C Action Space Design Based on Convex Region
The convex obstacle-free region is generated by the algorithm proposed by Zhong et al. [26], which can efficiently create reliable convex spaces among obstacles of any shape from LiDAR point cloud data.
To achieve efficient dynamic obstacle avoidance, the agent’s action is designed as a sequence of reference points at specific time intervals. In order to simplify the training process, this sequence is simplified to two points: one that represents the robot’s reference position after a control cycle ($t_{c}$) and another that represents after an MPC prediction horizon ($T$). These points serve as the short-term and long-term reference points, respectively. The path from the robot’s current position to these points forms the reference trajectory for MPC tracking.
When designing the action space, we first ensure that the sampling of reference points is confined within the convex region. Then, centering on the robot’s current position $O$ and considering its kinematic limits, we define two circular areas. These circles, intersected with the convex region, represent our short-term and long-term action spaces, as illustrated by the red areas in Fig. 2. The policy network is designed to output two sets of two-dimensional data. These data, after sigmoid activation, yield $\hat{a}_{t}=\left\{\left(\alpha_{t}^{s},\beta_{t}^{s}\right),\left(\alpha_{t}^{l},\beta_{t}^{l}\right)|\alpha_{t},\beta_{t}\in\left(0,1\right)\right\}$.
Figure 2: Schematic of Short-Term and Long-Term Reference Point Formulation
The short-term and long-term reference points coordinates $Q_{t}^{s}$ and $Q_{t}^{l}$, are calculated as Equation 2, with Fig. 2 illustrating this process. A ray at angle $\theta_{t}^{s}$ intersects with the short-term action space at $P_{t}^{s}$. We determine $P_{t}^{s}$ by calculating the polar angles of the convex’s vertices and finding which edge the angle $\theta_{t}^{s}$ falls within. This intersection, is solved by combining the ray and the edge equations, yields $P_{t}^{s}$. The distance $d_{\theta,t}^{s}$, from $O$ to $P_{t}^{s}$, is multiplied by the scaling factor $\beta_{t}^{s}$ to yield the distance $d_{t}^{s}$. Then we establish a new polar coordinate system centered at the short-term reference point $Q_{t}^{s}$. The long-term reference angle $\theta_{t}^{l}$, is obtained by adding the increment angle $\alpha_{t}^{l}\cdot 2\pi$ to $\theta_{t}^{s}$. Following the procedure used for the short-term point, we can determine the coordinates of $Q_{t}^{l}$.
@@
$\theta_{t}^{l}$ $=\theta_{t}^{s}+\alpha_{t}^{l}\cdot 2\pi$
@@
To navigate safely in unknown environment, our method utilizes real-time LiDAR data to construct the convex obstacle-free region for local navigation. By keeping the reference trajectory within the latest convex region, we can greatly reduce the risk of colliding with obstacles. The reference trajectory, connecting both long-term and short-term reference points is depicted by the green solid line in Fig. 3. Subsequently, the MPC framework solves the local navigation problem by calculating optimal control inputs so that the robot can continuously follow the reference trajectory. This process is iterated, ensuring the robot can follow the updated trajectory and reach the target without colliding with obstacles.
@@ III-D Model Predictive Control Formulation
After discretizing the integral model that governs robotic dynamics (as detailed in Section III-B), we derive a sequential relationship expressed as $x_{i}=Fx_{i-1}+Gu_{i-1}$ for $i=\{1,\ldots,N\}$. In this context, $F$ represents the matrix for discrete-time state transitions, and $G$ refers to the matrix for discrete-time control inputs.
Using the method mentioned in Section III-C, we get the convex region from point cloud data, with its vertices $P_{j}^{c},j=\left\{1,...,rnum_{v}\right\}$ arranged in clockwise direction. This region constrains the MPC-optimized points to enhance safety by reducing the risk of collisions with obstacles. This constraint is efficiently met by determining whether the trajectory point $Q_{i}$ is inside the convex region through the vector cross product method $\overrightarrow{Q_{i}P_{j}^{c}}\times\overrightarrow{Q_{i}P_{j+1}^{c}}\leq 0$.
During navigation, this vector method is also used to verify if the goal endpoint lies within the convex region. If so, this endpoint is directly used as the long-term reference point $Q_{t}^{l}$ , which promotes early training by facilitating the acquisition of high-reward trajectories. Additionally, we introduce a final state stop cost $\mathrm{cost}_{stop}$, defined in Equation 3, into the MPC’s cost function when the goal is within the convex region. This cost aims to minimize both velocity $V_{N}=\left(v_{x}^{N},v_{y}^{N}\right)$ and acceleration $A_{N}=\left(a_{x}^{N},a_{y}^{N}\right)$ at the destination, ensuring a controlled and safe termination of movement.
The short-term and long-term reference points are used to guide the positions of the first and last control period states predicted by the MPC. The cost function incorporates discrepancies between MPC-predicted points and reference points to ensure fidelity to the reference trajectory. Ultimately, the MPC is defined as a quadratic programming problem, formulated in Equation 4, optimizing the control sequence $u_{0:N-1}^{*}$. The optimization process generates a sequence of MPC-optimized points $Q_{i}^{*},i=\left\{1,...,N\right\}$, forming the MPC-optimized trajectory, as depicted in Fig. 4.
$\mathrm{cost}_{stop}=\begin{cases}\begin{aligned} &w_{vend}\|V_{N}\|^{2}\\ &+w_{aend}\|A_{N}\|^{2}\end{aligned}&\text{if goal in convex}\\ 0&\text{otherwise}\end{cases}$ (3)
@@
$\quad+w_{smooth}\sum_{k=0}^{N-1}{\left\|u_{k}\right\|^{2}}+\mathrm{cost}_{stop}$